A 50 g rock is launched vertically in the air by a slingshot. The slingshot exerts an average force of 53 N on the rock over a distance of 0.750 m.  


  1. Find the maximum height reached by the rock in [m].
  2. Find the speed of the rock as it leaves the slingshot in [m/s]?
1
Expert's answer
2020-08-03T15:11:04-0400

We can calculate the total energy given by the slingshot to the rock in form of "E = \bar{F}\cdot S = 53\,\mathrm{N}\cdot 0.750\,\mathrm{m} = 39.75\,\mathrm{J}."


2. This energy transforms into the kinetic energy of the rock, so the initial velocity can be calculated as "v = \sqrt{\dfrac{2E}{m}} = 39.9\,\mathrm{m/s}."


1. The rock stops to rise when the kinetic energy turns to 0. So the final potential energy is equal to the initial kinetic energy. Therefore,

"mgh = \dfrac{mv^2}{2} = E, \;\; h = \dfrac{E}{mg} \approx 81\,\mathrm{m}."